OpenStudy (anonymous):

2x+2/3z-8y x=7.2, y=3/4, z=3?

OpenStudy (campbell_st):

1st change y from 3/4 to 0.75 you need to substitute and evaluate \[\frac{(2 \times 7.2)+2}{(3 \times 0.75) -(8 \times 3)}\]

OpenStudy (anonymous):

Thank you very much. I didnt know how to do that. Thanks again.